Write the equation of a line that passes through (-1, 2) and (3, 5). Write your answer in slope-intercept form.
Answer by Alan3354(69443)

Write the equation of a line that passes through (-1, 2) and (3, 5). Write your answer in slope-intercept form.
--------
Find the slope m
m = diffy/diffx = 3/4
----
Use y = mx + b and either point to find b, the y-intercept
Using (3,5)
5 = (3/4)*3 + b
b = 5 - 9/4 = 7/4
-----
y = (3/4)x + 7/4